Determine the power series solution to the equation d2y/dx2+6xdy/dx-4y=0

about the point x0=0


1
Expert's answer
2021-08-09T16:13:22-0400

Given equation


"\dfrac{d^{2}y}{dx^{2}}+6x\dfrac{dy}{dx}-4y=0" about the point "x_{0}=0"


Let the solution of given series will be


"y=a_0x^{m}+a_1x^{m+1}+a_2x^{m+2}+a_3x^{m+3}+...."


"\dfrac{dy}{dx}=ma_0x^{m-1}+a_1(m+2)x^{m+1}+(m+3)a_2x^{m+2}+....."


"\dfrac{d^{2}y}{dx^{2}}=m(m-1)a_0x^{m-2}+a_1(m+2)(m+1)x^{m}+(m+3)a_2x^{m+1}+...."



Now putting , in the above equation we get -


"m(m-1)a_0x^{m-2}+a_1(m+2)(m+1)x^{m}+(m+3)a_3x^{m+1}+....6x[( ma_0x^{m-1}+a_1(m+2)x^{m+1}+(m+3)a_2x^{m+2}+.....)]-4[(a_0x^{m}+a_1x^{m+1}+a_2x^{m+2}+a_3x^{m+3}+....)]=0"

"a_0x^{m-2}[m(m-1)]+x^{m}[a_1(m+2)(m+1)+6ma_0-4a_0]+x^{m+1}[a_3(m+3)-4a_1]+x^{m+2}[6a_1(m+2)-4a_2]+x^{m+3}[6(m+3)a_2-4a_3]=0"



On equating the coefficient the lower power of "x," we get


The lowest power of "x" is "x^{m-2}" ","


On equation the lowest power of "x" which is "x^{m-2}" on both side of equation we get ,


"a_0m(m-1)=0" , which gives

"a_0=0" "m=0,1"


The solution of indicial equation , is given as


"\therefore y=c_1y_1+c_2y_2"


Now , on equating the coefficient of "x^{m}" , we get


"=" "[a_1(m+2)(m+1)+6ma_0-4a_0]" "=0"

"=" "[a_1(m+2)(m+1)+a_0(6m-4)]=0"


"a_1=-\dfrac{a_0(6m-4)}{(m+2)(m+1)}"


Now , on equation the coefficient of "x^{m+1}," we get


"a_3(m+3)-4a_1=0"


"a_3=\dfrac{4a_1}{m+3}"


"a_3=-\dfrac{4}{(m+3)}\dfrac{a_0(6m-4)}{(m+2)(m+1)}"


Now , on equating the coefficient of "x^{m+2}," we get


"6a_1(m+2)-4a_2=0"


"a_2=\dfrac{6a_1(m+2)}{4}"


"a_2=\dfrac{6(m+2)}{4}\dfrac{a_0(6m-4)}{(m+2)(m+1)}"


Thus , for "m=0" , we get


"y_{(m=0)}=y_1=[x^{m}(a_0+a_1x+a_2x^{2}+.....)]_{m=0}"


"=a_0-2a_0x-6a_0x^{2}+...."


"y_{m=0}=a_0(1-2x-6x^{2}+...)"


"y_{m=1}=y_2=[x^{m}(a_0+a_1x+a_2x^{2}+...)"


"y_2=x(a_0-\dfrac{a_0}{3}x+\dfrac{9}{2}\dfrac{a_0}{3}x^{2}+....)"



"y_2=a_0x(1-\dfrac{a_0}{3}x+\dfrac{3}{2}a_0x^{2}+...)"


The solution of given , equation is


"y=c_1a_0(1-2x-6x^{2})+c_2a_0x((1-\dfrac{a_0}{3}x+\dfrac{3}{2}a_0x^{2}+...)"